Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Difficult Math Question #9

Expert replies
by 800guy » Thu Sep 07, 2006 2:26 pm
i'll post the OA when a few people have responded...

A bag contains 3 red, 4 black and 2 white balls. What is the probability of drawing a red and a white ball in two successive draws, each ball being put back after it is drawn?

(A) 2/27
(B) 1/9
(C) 1/3
(D) 4/27
(E) 2/9
Join the discussion
Source: — Problem Solving |

by anupam » Thu Sep 07, 2006 9:40 pm
P(R)*P(W) * 2

4/27
Join the discussion

by Rishabh » Fri Sep 08, 2006 6:20 am
Soln:
3 red, 4 black and 2 white

total 9 balls

the first one can be either a red or a white
thus p(R1) = 3/9
and p(W1) = 2/9

since the balls are replaced, the second one can be either a white or red
thus p(W2) = 2/9
and p(R2) = 3/9

thus probability to draw either a red or white ball
= p(R1)*p(W2) + p(W1)*p(R2)
= 3/9*2/9 + 2/9*3/9
= 2(2/27)
= 4/27

Ans: D
Join the discussion

Re: Difficult Math Question #9

by piren » Sun Sep 10, 2006 6:04 am
800guy wrote:i'll post the OA when a few people have responded...

A bag contains 3 red, 4 black and 2 white balls. What is the probability of drawing a red and a white ball in two successive draws, each ball being put back after it is drawn?

(A) 2/27
(B) 1/9
(C) 1/3
(D) 4/27
(E) 2/9
(probability of a white and a red, whatever the order):

Hence
P(take a red and then a white) = 3/9 * 2/9 = 2/27
P(take a white and then a red) = 2/9 * 3/9 = 2/27
=> Total P =4/27, answer D.

IMHO, not a very difficult probability problem... I've seen much worse
Join the discussion

OA

by 800guy » Mon Sep 11, 2006 7:37 pm
OA:

Case I: Red ball first and then white ball

P1 = 3/9*2/9= 2/27

Case 2: White ball first and then red ball

P2 = 2/9*3/9 = 2/27

Therefore total probability: p1 + p2 = 4/27
Join the discussion

by naveen1003 » Thu Jun 09, 2011 3:08 am
A bag contains 3 red, 4 black and 2 white balls. What is the probability of drawing a red and a white ball in two successive draws, each ball being put back after it is drawn?

(A) 2/27
(B) 1/9
(C) 1/3
(D) 4/27
(E) 2/9
3R, 4B, 2W balls. Total = 9 balls
prob(RW) = Prob(R)*Prob(W) (independent events as the ball is put back after drawn)
= (3/9)*(2/9) = 2/27
Also the case when White ball is drawn first and Red ball next should also be considered.
prob(WR) = Prob(W)*Prob(R) = 2/27
prob(drawing red and white ball in 2 successive draws) = 2/27 + 2/27 = 4/27
IMO D
Join the discussion